solve the equation/show work please
square root of 3x -4 =2
___
5

Answer by checkley75(3666)

IF THIS IS
3X/5-4=2 THEN:
3X/5=2+4
3X/5=6 NOW CROSS MULTIPLY:
3X=5*6
X=30/3
X=10 ANSWER.
PROOF
3*10/5-4=2
30/5-4=2
6-4=2
2=2